3. Gains from trade Consider two neighboring island countries called Felicidad and Arcadia. They each have 4 million labor hoIn the following table, enter each countrys production decision on the third row of the table (marked Production). Suppose
Answer #1

Felicidad's opportunity cost of producing 1 pair of jeans is 32/8 = 4 bushel of rye and Arcadia's opportunity cost of producing 1 pair of jeans is 24/12 =2 bushel of rye. Because Arcadia has a lower opportunity cost for the production of jeans. Therefore, Arcadia has a comparative advantage in the production of jeans and Felicidad has a comparative advantage in the production of rye.

Suppose that each country completely specializes in the production of the good in which it has a comparative advantage , producing only that good . In this case,the country that produces jeans i.e Arcadia will produce (4)(12) = 48 million pairs per week And the country that produces rye i.e Felicidad will produce 4(32)=128 million bushels per week.

Suppose the country that produces corn trades 26 million pairs of jeans to the other country in exchange for 78 million bushels of rye.

FELICIDAD ARCADIA
Jeans (millions of pairs) Rye (Millions of bushels) Jeans (Millions of pairs) Rye (Millions of bushels)
Without trade Production 24 32 12 72
Consumption 24 32 12 72
With Trade Production 0 128 48 0
Trade action Import 26 Export 78 Export 26 Import 78
Consumption 26 (128-78)= 50 (48-26)=22 78
Gains from trade Increase in consumption (26-24)=2 (50-32)= 18 (22-12)=10 (78-72)=6

When the two countries did not specialize ,the total production of jeans was 36 million pairs per week and the total production of rye was 104 millions bushels per week. Because of specialization ,the total production of jeans was 48 million pairs per week and total production of rye was 128 million bushels per week.This implies that because of specialization , total production of jeans has increased by (48-36)=12 million pairs per week and total production of rye has increased by (128-104)=24 million bushels per week.